West Corp. issued 25-year bonds two years ago at a coupon rate of 5.3 percent. The bonds make semiannual payments. If these bonds currently sell for 105 percent of par value, what is the YTM?

DTO, Inc., has sales of $41 million, total assets of $27 million, and total debt of $8 million.


a. If the profit margin is 7 percent, what is the net income?


b. What is the ROA?


c. What is the ROE?
Returns
Year X Y
1 12 % 25 %
2 28 34
3 9 13
4 7 27
5 10 14

  

Using the returns shown above, calculate the arithmetic average returns, the variances, and the standard deviations for X and Y. (Do not round intermediate calculations. Enter your average return and standard deviation as a percent rounded to 2 decimal places, e.g., 32.16, and round the variance to 5 decimal places, e.g., .16161.)
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =23x2
1050 =∑ [(5.3*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^23x2
                   k=1
YTM% = 4.93

A

Profit margin% = Net income/sales
0.07 = Net income/41
Net income = 2.87
B
ROA = Net income/total Asset
ROA% = 2.87/27
ROA% = 10.63

C

Equity = assets-liabilites = 27-9 = 19

ROE = Net income/total equity
ROE% = 2.87/19
ROE% = 15.11

Please ask remaining parts separately, questions are unrelated. I have done one bonus